OpenStudy (anonymous):

Which of the following is a correct Pythagorean Theorem formula for the right triangle given below?

OpenStudy (anonymous):

t2 - p2 = d2 t2 + p2 = d2 p2 + d2 = t2 p2 - d2 = t2

OpenStudy (dumbcow):

answer is t^2+p^2=d^2 sum of squares of two legs = square of diagonal